Epidemiology question


3. ng the screening cut-off point affect sensitivity. specificity, PPV and NPV of the test? Why? are the respective costs ring the screening cut-off point affect the number of false positives and false negatives associated with false positives and false negativesi ? What 5. sensitivity, PPV and NPV of the test? How would you expect raising the cut-off point to 220mg of blood glucose per 100ml to affect the specificity was repeated using the 130mg of blood glucose per 100ml cut -off point in a population with an Calculate the sensitivity, specificity, 8% prevalence of diabetes, PPV, and NPV. How question 2? How does the increased prevalence of diabetes affect the PPV? as contrasted to the 5% prevalence in question 2, above. do the sensitivity and specificity compare to what they were in the 5% diabetic population in Gold-Standard Test No Disease Total 64 Disease 263 537 800 Screening Positive 327 9673 10000 9136 Negative 9200 Total Sensitivity Specificity PPV NPV
Answer #1

1. If the cut off point is less greater is the sensitivity and lowers specificity so more false positives.

If the cut off pointincreases then it lowers the sensitivity and highers the specificity gives more false negatives.

...................................................................................

If the glucose cut off point is made 220 /100 ml then true negatives then more false negatives will occur. People will have the disease but it will not be detected.

Ans6:

Sensityvity= a/a+c*100=263/800*100=32.875

Specificity=d/d+b*100=9163/9200*100=99.597

PPV=a/a+b*100=263/327*100=80.428

NPV=d/d+c*100=9163/9673*100=94.45